OpenStudy (anonymous):

Help! A spherical balloon is to be deflated so that its radius decreases at a constant rate of 7 cm/min. At what rate must air be removed when the radius is 5 cm? Air must be removed at cm^3/min.

OpenStudy (anonymous):

@AuroraB

OpenStudy (anonymous):

@amistre64 help

OpenStudy (amistre64):

what formulas do you think we will need to work this with?

OpenStudy (anonymous):

cilinder?

OpenStudy (anonymous):

im sorry i just dont know

OpenStudy (anonymous):

like i was doing one then i did the derivative

OpenStudy (amistre64):

the word cylindar is used nowhere in the problem ... maybe try the formula for the volume of a shpere

OpenStudy (anonymous):

ohh ok

OpenStudy (amistre64):

so, what is that formula? we will need to do an implicit derivative on it and plug in the given information

OpenStudy (anonymous):

so from the that equation, whats the next step

OpenStudy (amistre64):

take the implicit derivative and fill in the pertinent information

OpenStudy (anonymous):

but like how would the equation look with the numbers

OpenStudy (anonymous):

i can do the implicit derivative after that

OpenStudy (amistre64):

i need you to type out the formula for the volume of a sphere ...

OpenStudy (amistre64):

the numbers are after the implicit derivative is taken :)

OpenStudy (anonymous):

V= (4/3)pi r^3

OpenStudy (amistre64):

good, now the derivative with respect to time is?

OpenStudy (anonymous):

you want me to take the derivative of that equation?

OpenStudy (amistre64):

of course

OpenStudy (amistre64):

the only way i know of to get a rate of change for volume is to take the derivative of volume ....

OpenStudy (anonymous):

4 pi r^2

OpenStudy (anonymous):

whats next

OpenStudy (amistre64):

almost .... the chain rule applies since this is not a derivative with respect to the radius, but with respect to time; so an r' pops out V = (4/3)pi r^3 V' = 4 pi r^2 r' now, they give us r and r' stated in the problem, so we just fill them in to solve for V'

OpenStudy (anonymous):

so how does it look

OpenStudy (amistre64):

how does what look?

OpenStudy (anonymous):

4pi(7^2)(5)

OpenStudy (anonymous):

like that?

OpenStudy (amistre64):

the radius is given as: r = 5 the rate of change of the radius is given as: r' = 7 so you have placed them in the wrong spots is all

OpenStudy (anonymous):

and after that i will get the answer right?

OpenStudy (amistre64):

yes :)

OpenStudy (anonymous):

and that will be in cm^3/min?

OpenStudy (amistre64):

yes